OpenStudy (anonymous):

What are the possible values of x if the distance between the points (x, 8) and (–5, 3) is units? {–2, 9} {9, 1} {–9, –1} {–8, –3}

OpenStudy (anonymous):

if the distance between the two points is radical41units

OpenStudy (phi):

do you know the distance formula to find the distance between 2 points?

OpenStudy (phi):

Here it is http://www.purplemath.com/modules/distform.htm

OpenStudy (phi):

the first step is write down the distance formula. can you do that?

OpenStudy (anonymous):

yes i've done that

OpenStudy (phi):

second step is substitute (x,8) for (x1,y1) in the formula, and (-5,3) for (x2,y2) in the formula what do you get?

OpenStudy (anonymous):

\[d=\sqrt{(-5-x)^2 + (3-8)^2}\]

OpenStudy (phi):

good. also replace d with sqrt(41) now square both sides.

OpenStudy (anonymous):

1680=(-5-x) + (3-8) ?

OpenStudy (phi):

lost the bubble. first 41^2= 41*41= 1681 (not 1680) however, the equation you start with is \[ \sqrt{41}= \sqrt{ (-5-x)^2+ (3-8)^2} \] if we square both sides, what do we get note: sqrt(41)*sqrt(41) is not 1681.

OpenStudy (anonymous):

41=(-5-x) + (3-8)

OpenStudy (phi):

left side looks good. but try again on the right side when you multiply sqrt(a)*sqrt(a) you get a if a is something complicated, it does not matter. the only thing that changes is that the sqrt sign goes away.

OpenStudy (anonymous):

confused about the left side :/

OpenStudy (phi):

you mean the right side? sqrt(a+b+3)*sqrt(a+b+3) = a+b+3 sqrt(x^2/7)*sqrt(x^2/7) = x^2/7 sqrt( x^2 + y^2)* sqrt(x^2 + y^2) = x^2+y^2 do you see the pattern. what do you get with sqrt( (-5-x)^2 + (3-8)^2)*sqrt( (-5-x)^2 + (3-8)^2)=

OpenStudy (anonymous):

sorry i did mean the right. and i see the pattern. and would you get (-5-x)^2 + ( 3-8)^2 ?

OpenStudy (phi):

yes, so now we have 41= (-5-x)^2 + (3-8)^2 next step. simplify (3-8)^2

OpenStudy (anonymous):

25?

OpenStudy (phi):

you should get (3-8)= -5 and -5*-5 = 25 now expand (-5-x)^2 : (-5-x)*(-5-x) use FOIL to multiply this out

OpenStudy (anonymous):

x^2+10x+25

OpenStudy (phi):

so the equation is now 41= x^2+10x+25 + 25 can you simplify and solve for x?

OpenStudy (anonymous):

wait how come there is an extra 25?

OpenStudy (anonymous):

never mind. i got it

OpenStudy (phi):

Your question is the reason it is always good to write down the original equation in its entirety, and as you simplify parts of it, rewrite the whole thing. otherwise you might leave something out.

OpenStudy (anonymous):

i got stuck at -9=x^2+10x

OpenStudy (phi):

add 9 to both sides to get -9+9= x^2+10x+9 or x^2+10x+9=0 can you factor this?

OpenStudy (phi):

one way is list the factors of 9 1,9 3,3 do any of the pairs add to 10? yes the first pair so (x+1)(x+9)=0 now find the 2 x values that make this 0

OpenStudy (phi):

(x+3)(x+3)= x^2+6x+9 close, but no cigar

OpenStudy (anonymous):

no cigar ? o.o

OpenStudy (phi):

Just an expression... when you have time, this video on factoring is useful http://www.khanacademy.org/math/algebra/polynomials/v/factoring-quadratic-expressions meanwhile, the answer here is (x+1)(x+9)=0 this means either (x+1)=0 or (x+9)+0 the first equation x+1=0 becomes x=-1 and the 2nd gives us x= -9 so -1 and -9 are the answers.

OpenStudy (anonymous):

alright. thank you